Hello,
My name is Kristin and my time for the college placement exam is approaching and I need help!!!!!so please help
f(a)=2a-1 and g(a)=a^2, then g(f(a))=?

I believe this is setup like a^2(2a-1), but then I get confused after this point. I don't know whether I should multiply them by each other and then pull out the a^2, which would bring back to the initial equation. I am lost.
THANK YOU IN ADVANCE,
KRISTIN

Answer by venugopalramana(3286)

f(a)=2a-1 and g(a)=a^2, then g(f(a))=?
I believe this is setup like a^2(2a-1), but then I get confused after this point. I don't know whether I should multiply them by each other and then pull out the a^2, which would bring back to the initial equation. I am lost.
EASIER WAY TO GO ABOUT IS PUT F(A)=T...SINCE WE HAVE TO FIND G(F(A)...SO WE HAVE TO FIND G(T)=T^2...............................I
NOW SINCE F(A)=T...WE GET F(A)=2A-1=T.....SO SUBSTITUTING IN EQN.I ABOVE,WE GET THE ANSWER
G(F(A))=G(T)=T^2=(2A-1)^2...IS THE ANSWER